# Exploring the Robustness and Transferability of Patch-Based Adversarial Attacks in Quantized Neural Networks

### v1

@@ 4.7 Experiment 6: Impact of Patch Size and Visibility
As shown in Table 4, 8x8 and 10x10 patches achieve the highest attack success rates across all bitwidths, with mean success rates around 89.92% at 32-bit and remaining above 73% at 2-bit. The 6x6 patch shows a slightly lower success rate than the 8x8 and 10x10 patches, indicating that moderate increases in patch size can enhance adversarial impact without a significant drop in effectiveness at lower bitwidths. The 12x12 patch, however, has a somewhat lower success rate, especially at higher bitwidths (32 and 8-bit), likely due to its larger size disrupting the natural distribution more noticeably, which may reduce its effectiveness.
Across all patch sizes, attack success rates generally decrease as bitwidth is reduced, with a consistent decline from 32-bit to 2-bit. Nevertheless, even at 2-bit, attack success rates remain relatively high (above 72% for most patch sizes). This modest decline suggests that while quantization reduces the patch attack’s effectiveness to some extent, it is insufficient to fully neutralize the threat. The high visibility and contrast of the patch likely contribute to its resilience across bitwidths.
The 8x8 and 10x10 patches exhibit the highest attack success rates, indicating that increasing patch size improves the attack’s effectiveness up to a certain point. This supports the hypothesis that larger, more visible patches exert a stronger adversarial influence that is less affected by quantization. The slight drop in effectiveness for the 12x12 patch may suggest that excessively large patches interfere with the model’s natural feature processing, thereby slightly diminishing the adversarial impact.
@@
6x6 86.55 78.47 75.81 84.53 72.52
8x8 89.92 81.94 75.22 81.16 73.37
10x10 89.92 81.94 75.22 81.16 73.37
12x12 84.87 75.00 73.41 76.27 78.51
Table 4: Attack Success Rates (%) for different patch sizes on Resnet-56.
@@ 6.1.1 Quantization in Neural Networks
6.1.2 Challenges of Attacking Quantized Neural Networks
Attacking quantized models poses unique challenges due to the discretization of weights and activations, which fundamentally alters the model’s gradient landscape. Two key factors, quantization shift and gradient misalignment, hinder the effectiveness and transferability of traditional adversarial examples on QNNs.
  1. 1.
    Quantization Shift: Quantization reduces continuous values to discrete buckets, causing clean and adversarial features to map into the same quantization levels. This phenomenon, referred to as quantization shift, can significantly weaken the adversarial effect. In traditional full-precision models, adversarial perturbations are designed to push the model’s features in a specific direction that leads to misclassification. However, in quantized models, both clean and perturbed features may be rounded to the same value, nullifying the adversarial intent.
  2. 2.
    Gradient Misalignment: To enable backpropagation in quantized models, quantization-aware training commonly employs the Straight-Through Estimator (STE) [2], an approximation technique that allows gradients to pass through discrete operations. Although effective for training, STE introduces inaccuracies in the gradient calculations, leading to gradient misalignment between full-precision and quantized models. This misalignment makes it difficult to transfer adversarial examples across models with different bit widths or quantization settings, as the gradient directions no longer align.
@@ 6.1.4 Patch-based Adversarial Attacks
  • Google Adversarial Patch (GAP): The Google Adversarial Patch (GAP) [4] emphasizes practicality in real-world attacks, addressing challenges associated with traditional $L_{p}$-norm-based adversarial perturbations, such as requiring precise object capture through a camera. GAP creates universal adversarial patches that can be applied to any part of an input, making it more versatile. Furthermore, it incorporates the Expectation over Transformation (EOT) [1] framework to improve the patch’s robustness against common variations like rotation, scaling, and translation, ensuring its effectiveness under real-world conditions.
@@ 6.3.2 Experiment 2: Effect of Quantization Levels on Patch Transferability
ResNet-20 mean 84.71 49.61 65.31 50.45 58.31
std 1.63 2.25 2.59 1.32 2.82
ResNet-56 mean 84.40 56.69 65.31 54.22 47.91
std 1.13 2.64 2.59 2.69 1.94
VGG-16 mean 95.79 39.65 63.76 48.7 40.69
std 0.51 0.93 2.04 2.01 1.83
VGG-19 mean 95.71 52.04 55.03 48.90 64.24
std 0.71 1.80 1.44 2.14 1.58
@@
6.3.3 Experiment 3: Spatial Sensitivity of Patches
In this part of the experiment, we assess the robustness of patch-based adversarial attacks to rotational transformations. Specifically, we evaluate the attack success rates when the patch is rotated at different angles ($0^{\circ}$, $5^{\circ}$, $10^{\circ}$, $20^{\circ}$, and $30^{\circ}$) and applied to ResNet-56 quantized to different bit widths (32-bit, 8-bit, 5-bit, 4-bit, and 2-bit).
The results (Table 6) indicate that patches maintain high attack success rates across all tested rotation angles, demonstrating resilience to spatial orientation changes. This robustness persists across different quantization levels, although success rates decline slightly at lower bit widths (e.g., 2-bit).
Angle 32-bit 8-bit 5-bit 4-bit 2-bit
Rotated $0^{\circ}$ 86.059 85.775 77.875 80.631 77.06
Rotated $5^{\circ}$ 87.361 85.987 79.268 85.811 78.842
Rotated $10^{\circ}$ 86.245 87.686 79.791 84.459 79.51
Rotated $20^{\circ}$ 84.387 84.501 77.178 80.856 81.737
Rotated $30^{\circ}$ 85.987 85.987 82.056 78.829 79.065
Table 6: Attack Success Rates (%) for different rotation angle on ResNet-56.

### v2

@@ 4.7 Experiment 6: Impact of Patch Size and Visibility
We generate patch-based attacks with varying patch sizes on a 32-bit model and we evaluate the success rate of each patch size for different quantization levels. As shown in Table 8, attack success rates generally decrease as bitwidth is reduced in all patch sizes, with a consistent decline from 32-bit to 2-bit. The 10x10 and 12x12 patches exhibit the highest attack success rates, indicating that increasing patch size improves the attack’s effectiveness. This supports the hypothesis that larger, more visible patches exert a stronger adversarial influence that is less affected by quantization.
@@
6x6 86.43 83.24 79.94 76.22 73.08
8x8 89.92 86.94 81.62 79.76 75.18
10x10 91.92 88.94 83.50 81.16 78.62
12x12 95.56 90.10 88.52 86.19 84.89
Table 8: ASR (%) of LAVAN attack against ResNet-56 on CIFAR-10 for different patch sizes.
@@ 5 Key Strategies for Defense Design
6 Quantization-Aware Defense Training (QADT)
Based on our findings, we introduce Quantization-Aware Defense Training (QADT), a strategy that integrates adversarial patch augmentation to improve robustness against patch-based attacks in quantized models. QADT trains models with adversarial patches crafted under simulated quantized conditions, enabling them to adapt to altered gradients and mitigate low-bitwidth vulnerabilities. Additionally, training with diverse patch variants (e.g., varying in size, position, and rotation) enhances spatial and feature-level awareness, reducing overfitting to specific configurations.
Input: Dataset $D$, model $M$, learning rate $\eta$, number of epochs $E$, batch size $B$, adversarial patch generator $P$, quantization levels $Q$.
Output: Robust quantized model $M_{QADT}$.
1. Initialize model $M$ with random weights.
2. for epoch $e=1,\dots,E$:
3:       Shuffle $D$ and split into batches of size $B$.
4:       for each batch $b\in D$:
5:           Extract clean samples $(x,y)\in b$.
6:           Generate diverse adversarial patches $p$ with generator $P$ for quantization levels $Q$:
7:               Vary patch size, position, rotation, and intensity.
8:               Apply simulated quantization on $p$ using levels $Q$.
9:           Create adversarial samples $(x_{adv},y)$ by embedding patches $p$ into $x$.
10:           Perform forward pass with $(x,y)$ and $(x_{adv},y)$:
11:               Compute loss $L=\text{Loss}(M(x),y)+\text{Loss}(M(x_{adv}),y)$.
12:           Perform backward pass and update $M$ with $\eta$.
13:       end for
14: end for
15: Quantize model $M$ to levels $Q$ for deployment.
16: return Quantized robust model $M_{QADT}$.
Algorithm 1 Quantization-Aware Defense Training (QADT)
The QADT algorithm 1 aims to train a robust neural network model against adversarial patch attacks while accounting for model quantization effects. The process begins with the initialization of the model $M$ with random weights. During training, the algorithm iterates over a specified number of epochs $E$, shuffling the dataset $D$ and splitting it into batches of size $B$ in each epoch. For each batch, clean samples $(x,y)$ are extracted, and a diverse set of adversarial patches $p$ is generated using the adversarial patch generator $P$, considering different quantization levels $Q$. The diversity of the patches is enhanced by varying size, position, rotation, and intensity, followed by simulated quantization to align with the target quantization levels. These adversarial patches are then embedded into the clean samples to produce adversarial inputs $(x_{adv},y)$. Both clean and adversarial samples undergo a forward pass through the model, with the total loss $L$ calculated as the sum of the classification losses for both sample types. The model’s weights are updated using backpropagation with a learning rate $\eta$. This process is repeated for all batches across all epochs. Upon completion of training, the model is quantized to the desired levels $Q$ to ensure it is optimized for deployment. The final output is a robust, quantized model $M_{QADT}$ that is resilient to adversarial patch attacks and well-adapted to the challenges of quantized architectures.
As shown in Table 9, QADT significantly reduces the attack success rate, achieving 21.3% ASR (32-bit) and 10.65% ASR (8-bit), a substantial improvement over standard training (ST) and adversarial training with PGD-7 (AT).
ST AT QADT ST AT QADT
Patch Size 32-bit 32-bit 32-bit 8-bit 8-bit 8-bit
(10x10) 87.29 44.01 21.3 57.51 39.69 10.65
Table 9: ASR (%) on different trained ResNet20 (CIFAR-10) using the LAVAN attack.
@@ 7.1.1 Quantization in Neural Networks
Training Details: For CIFAR-10 dataset, all models were trained with a batch size of 128 using Stochastic Gradient Descent (SGD) with a momentum of 0.9 and weight decay set to $1\times 10^{-4}$. Each model was trained for 50 epochs to ensure adequate convergence across quantization levels and architectures.
@@ 7.1.3 Patch-based Adversarial Attacks
  • Google Adversarial Patch (GAP): GAP attack [3] emphasizes practicality in real-world attacks, addressing challenges associated with traditional $L_{p}$-norm-based adversarial perturbations, such as requiring precise object capture through a camera. GAP creates universal adversarial patches that can be applied to any part of an input, making it more versatile. Furthermore, it incorporates the Expectation over Transformation (EOT) [1] framework to improve the patch’s robustness against common variations like rotation, scaling, and translation, ensuring its effectiveness under real-world conditions.
  • Deformable Patch Representation (DPR): DRP attack [5] is a type of adversarial attack that focuses on altering the shape and appearance of patches in an image rather than just modifying pixel values. The main idea is to create deformable and dynamic patches that can adapt their form to exploit the vulnerabilities of neural networks more effectively. DPR essentially aims to increase the robustness and transferability of adversarial patches by introducing shape-based deformations, which makes the attack less dependent on specific pixel perturbations and more adaptable across different models and input variations.
@@ 7.3.2 Experiment 2: Effect of Quantization Levels on Patch Transferability
ResNet-20 84.71 49.61 65.31 50.45 58.31
ResNet-56 84.40 56.69 65.31 54.22 47.91
VGG-16 95.79 39.65 63.76 48.7 40.69
VGG-19 95.71 52.04 55.03 48.90 64.24